OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 2x – 3y = –29 x + 4y = 13

OpenStudy (anonymous):

a. 2x – 3(4y + 13) = –29 b. 2(4y + 13) – 3y = –29 c. 2(–4y + 13) – 3y = –29 d. 2x – 3(–4y + 13) = –29